OpenStudy (anonymous):

The length of an equilateral triangle is the same as the length of a rectangle and the width of the rectangle is 2 inches less than its length. If the perimeter of the triangle is 4 inches less than the perimeter of the rectangle, what are the dimensions of the rectangle?

OpenStudy (anonymous):

I would like to do it using algebra

OpenStudy (ajprincess):

Let the length of the triangle be x. So the perimeter of the triangle=3x Let the width of the rectangle=y y=x-2 Perimeter of the rectangle=2(x+y) =2(x+x-2) =2(2x-2) =4x-4 3x=4x-8 Can u find x nw?

OpenStudy (anonymous):

3x=4x-8 +8 Then it would become: 8+3x=4x -3x -3x Then it would be: 8=x

OpenStudy (anonymous):

Is that correct?

OpenStudy (ajprincess):

yeah u r right.:) Nw find y.

OpenStudy (ajprincess):

Can u?

OpenStudy (anonymous):

I would have to use y=x-2?

OpenStudy (anonymous):

right?

OpenStudy (ajprincess):

ya

OpenStudy (anonymous):

y=x-2 -x=-x So it would be: -x+y=-2 I don't think that is right :/

OpenStudy (ajprincess):

y=8-2 nw can u?

OpenStudy (anonymous):

subtract 8-2=6 y=6

OpenStudy (anonymous):

I'm sure that's right

OpenStudy (ajprincess):

nw u hav the dimensions of the rectangle.

OpenStudy (anonymous):

Thank you so much.

OpenStudy (ajprincess):

u r welcome.